# Interaction induced Anderson transition in a kicked one dimensional Bose gas

### v1

@@ Abstract
where $c=g/(\varepsilon\ell)$, $K=\mathcal{K}/(\tau\varepsilon)$, $t=t^{\prime}/\tau$, $x_{i}=y_{i}/\ell$, $p_{i}=-i\hbar_{e}\partial/\partial x_{i}$ the momenta of the particles satisfying the commutation relation $[x_{j},p_{j}]=i\hbar_{e}$, where $\hbar_{e}=\hbar/(\varepsilon\tau)$ is the effective Planck’s constant. $H_{0}$ is the Lieb-Liniger Hamiltonian which can be exactly solved with Bethe Ansatz LiebLiniger . The eigenstates are labeled by a set of $N$ rapidities $\lambda_{i}$ and can be written, in the fundamental sector $x_{1}\leq x_{2}\leq...\leq x_{N}$, as
@@
where the Bethe numbers $I_{j}$ are integers for odd $N$ and half-integers for even $N$, and must all be distinct. The energy of a Bethe state is given by $E_{\vec{\lambda}}=\frac{\hbar_{e}^{2}}{2}\sum_{i=1}^{N}\lambda_{i}^{2}$ and its momentum $P_{\vec{\lambda}}=\hbar_{e}\sum_{i=1}^{N}\lambda_{i}$. The ground state corresponds to the lowest available set of Bethe numbers. For instance, for $N=3$, it is given by $\vec{I}=(-1,0,+1)$. Due to the bosonic statistics, these states are symmetric under permutations of spatial coordinates and rapidities, allowing us to restrict our calculations to the fundamental rapidity sector $I_{1}<I_{2}<...<I_{N}$. It is important to note that the rapidities $\lambda_{i}$ differ from the actual momenta of the particles except in the non-interacting limit. In the limit of infinite interactions ($c\to+\infty$), the rapidities coincide with the momenta of the corresponding free fermions via the Bose-Fermi mapping Girardeau . In general, however, they do not reside on a regular lattice (which would have spacing $12\pi\ell/L=$ 1 in our units), but are instead coupled through the Bethe equations (5).
@@
and ${\rm Re}\{X\}$ denotes the real part of a complex number $X$. This equation closely resembles an Anderson model with long-range hopping in an $N$-dimensional rapidity space. However, for this analogy to hold rigorously, two key conditions must be met. First, the on-site energy $W_{\vec{\lambda},\vec{\lambda}}$ must exhibit pseudo-random behavior, which is ensured if $\hbar_{e}$ is incommensurate with $4\pi$ FishmanGrempelPrange . Second, the hopping terms must decay faster than $1/||\vec{\lambda}-\vec{\mu}||$ Mirlin1996 . This requirement is satisfied in our model since the matrix elements $M_{{\vec{\lambda}},{\vec{\mu}}}$ decay exponentially suppmat .
@@
Results for $N=3$ bosons.– To illustrate our findings, we consider the case of $N=3$ bosons, where extensive numerical calculations can be performed. To do so, we truncate the Hilbert space by restricting the Bethe numbers to the range $-N_{s}\leq I_{j}\leq N_{s}$, where $N_{s}$ is an integer (since $N=3$ is odd). The computation of the Floquet operator matrix elements (6) then reduces to evaluating $M_{{\vec{\lambda}},{\vec{\mu}}}=\langle{\vec{\lambda}}|e^{-iH_{K}/\hbar_{e}}|{\vec{\mu}}\rangle$ since $e^{-iH_{0}/\hbar_{e}}$ is diagonal in this basis. These matrix elements can be computed analytically using (3) suppmat . The total matrix size in the truncated basis is given by $N_{M}=\binom{2N_{s}+1}{3}$. We initialize the system in the ground state, characterized by Bethe numbers $\vec{I}=(-1,0,1)$, and evolve it numerically by iterating the Floquet operator in the truncated Hilbert space with $N_{s}=29$ for selected values of $K$, $c$, and $\hbar_{e}$. The total energy is then computed as $\langle E(t)\rangle=\sum_{\vec{\lambda}}|\alpha_{\vec{\lambda}}(t)|^{2}E_{\vec{\lambda}}$, where $\alpha_{\vec{\lambda}}(t)$ represents the amplitude of the many-body wave function in the Bethe state $|\vec{\lambda}\rangle$.
@@
Figure 3: (a) The rescaled quantity $\ln\Lambda=\ln[\langle E(t)\rangle t^{-2/3}]$ as a function of $K$ for different times between $t=10$ and $t=200$. All curves intersect approximately at the critical point $(K_{c}/\hbar_{e}\simeq 2.1,\ln\Lambda_{c}\simeq 1.54)$ demonstrating the existence of a metal-insulator transition. (b) Determination of the critical exponent $\nu$ by fitting $(\ln\Lambda)^{\prime}(K_{c},t)\sim t^{1/3\nu}$ in log-log scale. Parameters are the same as in Fig. 2.
@@ Acknowledgements.
We would like to acknowledge helpful discussions with L. Bellando De Castro, J. C. Garreau, B. Grémaud, F. Hébert, C. Miniatura, A. Rançon and F. Werner. This work has benefited from the financial support of Agence Nationale de la Recherche under Grant No ANR-21-CE47-0009 Quantum-SOPHA and Grant No ANR-23-PETQ-0001 Dyn1D France 2030.

### v2

@@ Abstract
where $c=g/(\varepsilon\ell)$, $K=\mathcal{K}/(\tau\varepsilon)$, $t=t^{\prime}/\tau$, $x_{i}=y_{i}/\ell$, $p_{i}=-i\hbar_{e}\partial/\partial x_{i}$ the momenta of the particles satisfying the commutation relation $x_{j},p_{j}]=i\hbar_{e}$, where $\hbar_{e}=\hbar/(\varepsilon\tau)$ is the effective Planck’s constant. $H_{0}$ is the Lieb-Liniger Hamiltonian which can be exactly solved with Bethe Ansatz LiebLiniger . The eigenstates are labeled by a set of $N$ rapidities $\lambda_{i}$ and can be written, in the fundamental sector $x_{1}\leq x_{2}\leq...\leq x_{N}$, as
@@
where the Bethe numbers $I_{j}$ are integers for odd $N$ and half-integers for even $N$, and must all be distinct. The energy of a Bethe state is given by $E_{\vec{\lambda}}=\frac{\hbar_{e}^{2}}{2}\sum_{i=1}^{N}\lambda_{i}^{2}$ and its momentum $P_{\vec{\lambda}}=\hbar_{e}\sum_{i=1}^{N}\lambda_{i}$. The ground state corresponds to the lowest available set of Bethe numbers. For instance, for $N=3$, it is given by $\vec{I}=(-1,0,+1)$. Due to the bosonic statistics, these states are symmetric under permutations of spatial coordinates and rapidities, allowing us to restrict our calculations to the fundamental rapidity sector $I_{1}<I_{2}<...<I_{N}$. It is important to note that the rapidities $\lambda_{i}$ differ from the actual momenta of the particles except in the non-interacting limit. In the limit of infinite interactions ($c\to+\infty$), the rapidities coincide with the momenta of the corresponding free fermions via the Bose-Fermi mapping Girardeau . In general, however, they do not reside on a regular lattice (which would have spacing $2\pi\ell/L=1$ in our units), but are instead coupled through the Bethe equations (5).
@@
[…]
@@
Results for $N=3$ bosons.– To illustrate our findings, we consider the case of $N=3$ bosons, where extensive numerical calculations can be performed. To do so, we truncate the Hilbert space by restricting the Bethe numbers to the range $N_{s}\leq I_{j}\leq N_{s}$, where $N_{s}$ is an integer (since $N=3$ is odd). The computation of the Floquet operator matrix elements (6) then reduces to evaluating $M_{{\vec{\lambda}},{\vec{\mu}}}=\langle{\vec{\lambda}}|e^{-iH_{K}/\hbar_{e}}|{\vec{\mu}}\rangle$ since $e^{-iH_{0}/\hbar_{e}}$ is diagonal in this basis. These matrix elements can be computed analytically using (3) suppmat . The total matrix size in the truncated basis is given by $N_{M}=\binom{2N_{s}+1}{3}$. We initialize the system in the ground state, characterized by Bethe numbers $\vec{I}=(-1,0,1)$, and evolve it numerically by iterating the Floquet operator in the truncated Hilbert space with $N_{s}=29$ for selected values of $K$, $c$, and $\hbar_{e}$. The total energy is then computed as $\langle E(t)\rangle=\sum_{\vec{\lambda}}|\alpha_{\vec{\lambda}}(t)|^{2}E_{\vec{\lambda}}$, where $\alpha_{\vec{\lambda}}(t)$ represents the amplitude of the many-body wave function in the Bethe state $|\vec{\lambda}\rangle$.
@@
Figure 3: (a) The rescaled quantity $\ln\Lambda=\ln[\langle E(t)\rangle t^{-2/3}]$ as a function of $K$ for different times between $t=5$ (light blue) and $t=150$ (dark blue). All curves intersect approximately at the critical point $(K_{c}/\hbar_{e}\simeq 2.1,\ln\Lambda_{c}\simeq 1.54)$ demonstrating the existence of a metal-insulator transition. (b) Determination of the critical exponent $\nu$ by fitting $(\ln\Lambda)^{\prime}(K_{c},t)\sim t^{1/3\nu}$ in log-log scale. Parameters are the same as in Fig. 2.
@@ Acknowledgements.
We would like to acknowledge helpful discussions with L. Bellando De Castro, J. C. Garreau, B. Grémaud, F. Hébert, M. Landini, C. Miniatura, A. Rançon and F. Werner. This work has benefited from the financial support of Agence Nationale de la Recherche under Grant No ANR-21-CE47-0009 Quantum-SOPHA and Grant No ANR-23-PETQ-0001 Dyn1D France 2030. Note.—During the publication process, we found that A. Yang et al Landini2025 independently posted a preprint on the same topic at the same time.